OpenStudy (tinylittlehelper):

Help please! Kareena has $20. She woke up and went shopping. She spent money on 5 apples, and each is $2. She drove to work and payed for 3 gallons on 46 cent a gallon. On the way, she found a $5 bill on the ground. When Kareena arrived, she worked PART of her night shift, which consists of $10 an hour and had 8 hours in all. Then she worked over time for 3 hours and earned 6/47 of what she earned on her night shift. She worked for 8 hours in total. She paid for dinner for $4 and 7.5% tax included. At the end of her day, how much money did she have?

OpenStudy (anonymous):

starts with 20 bought 5 apples at $2 per apple = $10 20 - 10 = 10....now she has $10 drove to work...bought gas...3 gallons at 0.46/gal = 3 * 0.46 = $ 1.38 10 - 1.38 = 8.62....current balance found $5....new balance = 8.62 + 5 = 13.62 worked...$10/hr for 8 hrs = 10 * 8 = 80 80 + 13.62 = 93.62 --- new balance I am stuck here...hold on

OpenStudy (anonymous):

@Kainui ...you got a second to look at this

OpenStudy (anonymous):

do I do 6/47 * 10 to figure how much she made per hr overtime or do I multiply 6/47 * 80 ?

OpenStudy (anonymous):

and then it says she worked for 8 hrs....at how much per hr ?

OpenStudy (anonymous):

the wording is confusing me...did she work a TOTAL of 8 hrs...of which 3 are overtime and 5 is regular pay at $10/hr ???

OpenStudy (anonymous):

@shamim ...help,please

OpenStudy (anonymous):

sorry ....I cannot finish